OpenStudy (anonymous):

In a card game,hearts and spades are assigned diffrent point values.Sheila has 3 hearts and 4 spades,which is worh 41 points.Diego has 6 hearts and 5 spades,which is worth 67 points.How many points is a heart worth???

OpenStudy (unklerhaukus):

S=3hearts+4spades=41p D=6hearts+5spades=67p

OpenStudy (unklerhaukus):

3h+4s=42 6h+5s=67

OpenStudy (unklerhaukus):

solve these equations simultaneously ,

OpenStudy (anonymous):

how do i do that..

OpenStudy (anonymous):

Wait hold on why is it 42 not 41

OpenStudy (unklerhaukus):

ops my error

OpenStudy (unklerhaukus):

3h+4s=41 imagine we had this hand, Twice, 6h+8s, which would be worth 82 points

OpenStudy (anonymous):

it would be the 3h+4s=41

OpenStudy (anonymous):

righhht

OpenStudy (unklerhaukus):

now we know 6h+8s=82 and we already knew diago's hand 6h+5s, which was worth =67 the only difference is three spades and (82-67) points

OpenStudy (unklerhaukus):

so we can determine the value of a single spade by dividing by three

OpenStudy (anonymous):

so we divide 82/3????

OpenStudy (unklerhaukus):

oh, dont divide by three yet find the difference first 82-67=?

OpenStudy (anonymous):

15

OpenStudy (unklerhaukus):

ok so that means 3spades=15p now divide both sides by three

OpenStudy (unklerhaukus):

1spade=?p

OpenStudy (anonymous):

so the answer would be 6

OpenStudy (anonymous):

1 spade = 1p

OpenStudy (unklerhaukus):

that is not right

OpenStudy (unklerhaukus):

1 spade =15/3 points

OpenStudy (anonymous):

okay.....

OpenStudy (unklerhaukus):

what is one spade worth is you simplify the quotient ?

OpenStudy (anonymous):

yeah im soo bad at math .

OpenStudy (unklerhaukus):

is 15 on your three times tables ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

3x5=15

OpenStudy (unklerhaukus):

good 3x5=15 so 5=15/3

OpenStudy (unklerhaukus):

right? so each spade is worth 5 points,

OpenStudy (anonymous):

yes

OpenStudy (unklerhaukus):

you can now uses this information when looking at Either of Sheila's 3hearts+4spades=41p Or Diago's hand 6hearts+5spades=67p

OpenStudy (anonymous):

So how do we find the points of heart

OpenStudy (unklerhaukus):

3h+4s=41p (1) 6h+5s=67p (2) if you multiply the both sides of equation(1) by 2 you get 6h+8s=82p (3) now what is the difference between (2) and (3) (2) - (3) = ?